OpenStudy (anonymous):

Simplify the following expression: (5x2 + 3x + 4) − (2x2 + 5x − 1). If the final answer is written in the form Ax2 + Bx + C, what is the value of A? Numerical Answers Expected! Answer for Blank 1:

OpenStudy (anonymous):

ok so do i do the same thing again? @AravindG

OpenStudy (aravindg):

yep :)

OpenStudy (aravindg):

add like terms !

OpenStudy (anonymous):

so 5x2+3x+2x2+5x and 4-1 are the like terms? @AravindG

OpenStudy (aravindg):

yep :) (5-2)x^2+(3-5)x+4-1 = ?

OpenStudy (aravindg):

oops its 4+1

OpenStudy (aravindg):

(5-2)x^2+(3-5)x+4+1 = ?

OpenStudy (anonymous):

15x2+5?

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

5x^2 - 2x^2 = 3x^2

OpenStudy (aravindg):

note you cannot add x^2 and x both are not like terms ..just do what i showed above you will have the asnwer

OpenStudy (anonymous):

@AravindG is right

OpenStudy (anonymous):

3x^2+-2x+5?

OpenStudy (anonymous):

correct :) but that isnt ur answer

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

@HawkCrimson what do you mean not my answer

OpenStudy (aravindg):

correct !! :)

OpenStudy (anonymous):

thank you guys!!

OpenStudy (anonymous):

they want to know the value of A now, so wat is the value of A according to ur equation?

OpenStudy (anonymous):

3x^2?

OpenStudy (anonymous):

the value of A is the coefficient of x^2 so ur answer is 3 :)

OpenStudy (anonymous):

A = 3

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

ur welcome :)

OpenStudy (aravindg):

:)